OpenStudy (anonymous):

The length of a rectangle is 5 meters more than its width. If the area is 66 square meters, what are the length and width?

OpenStudy (btaylor):

Area of a rectangle = L x W L = W + 5 66 = L x W 66 = W(W+5) W^2 + 5W = 66 W^2 + 5W - 66 = 0 (W+11)(W-6) = 0 W = -11 or W = 6 the answer can't be -11 because it is impossible for width to be negative. So, W=6 and L = 11